# An Ran Chen

### v1

@@ Abstract
Bug fixing is a crucial task in software maintenance to hold user trust. Although various automated fault localization techniques exist, they often require specific conditions to be effective. For example, Spectrum-Based Fault Localization (SBFL) techniques need at least one failing test to identify bugs, which may not always be available. Bug reports, particularly those with stack traces, provide detailed information on system execution failures and are invaluable for developers. This study focuses on utilizing stack traces from crash reports as fault-triggering tests for SBFL. Our findings indicate that only 3.33% of bugs have fault-triggering tests, limiting traditional SBFL efficiency. However, 98.3% of bugfix intentions align directly with exceptions in stack traces, and 78.3% of buggy methods are reachable within an average of 0.34 method calls, proving stack traces as a reliable source for locating bugs. We introduce a new approach, SBEST, that integrates stack trace data with test coverage to enhance fault localization. Our approach shows a significant improvement, increasing Mean Average Precision (MAP) by 32.22% and Mean Reciprocal Rank (MRR) by 17.43% over traditional stack trace ranking methods.
Keywords:
bug report, stack trace, empirical study, fault localization, software maintenance.
@@ 1 Introduction
Locating and fixing software bugs is an essential yet frequently expensive and time-consuming task. Bug-fixing activities not only delay the project’s progress but also increase the software’s overall cost, making them a critical part of software maintenance. Considering the complexity and rapid evolution of modern software systems, the ability to efficiently and accurately locate bugs is more important than ever. To address this challenge, previous research has proposed numerous different automated fault localization techniques (Abreu et al., 2007; Jones and Harrold, 2005; Naish et al., 2011; Abreu et al., 2009; Wong et al., 2013; Papadakis and Le Traon, 2015; Hong et al., 2017; Papadakis and Le Traon, 2012; Wang et al., 2022; Cui et al., 2020), among which Spectrum-Based Fault Localization (SBFL) approaches are prominent due to its high accuracy (de Souza et al., 2016). SBFL techniques leverage information derived from test executions — specifically, the spectrum of code elements executed during passing and failing tests — to calculate the probability of every code element being faulty (i.e. suspiciousness score).
The effectiveness of SBFL highly depends on the availability of failing test cases that uncover the bug (i.e., fault-triggering tests). However, studies indicate that in real-world scenarios, especially in continuous deployment and integration environments, failing tests may not always be present or identified when bugs are reported (Chen et al., 2023; Kabadi et al., 2023). This absence of fault-triggering tests significantly impacts the applicability of SBFL, as it depends on it to distinguish between faulty and non-faulty code. In the absence of failing tests, alternative approaches must be considered. Exception stack traces, commonly included in crash reports, provide dynamic execution information that can be extremely valuable for debugging. Studies show that developers usually rely on them when investigating bugs (Schroter et al., 2010; Zimmermann et al., 2010). In addition, previous research (Chen et al., 2021a) shows that stack traces provide valuable information on the cause of the bugs and their location in the source code. These stack traces, which outline the sequence of method calls leading up to an exception, offer implicit clues about where a bug may reside, providing a fallback localization method in environments where traditional test-based approaches falter.
@@
In this paper, we address the challenges of fault localization in scenarios where traditional fault-triggering test data is scarce. We conduct our study on the Defects4J 2.0 dataset (Just et al., 2014b), a widely recognized resource for research in fault localization and program repair. However, prior studies (Kabadi et al., 2023; Chen et al., 2023) have reported that some tests in Defects4J were added post hoc, following the reporting of bugs, which could bias the bug localization process. To avoid this issue and better simulate the state of the codebase at the time developers first address a bug, we utilize the commit data immediately preceding the bug report rather than relying directly on the buggy commits provided by Defects4J. Our initial analysis of the bugs reveals a critical limitation: only 3.33% of crash reports, which are bug reports that contain stack traces, possess fault-triggering tests. Even for bug reports without stack traces, only 10.19% of the bugs have fault-triggering tests. This low incidence of failing tests challenges the efficacy of Spectrum-Based Fault Localization (SBFL) methods, which rely heavily on test outcomes to identify bug locations. To overcome this limitation, we propose leveraging the stack traces—rich in contextual and execution information about the bug’s root causes—as a substitute for the failing tests. We introduce a novel approach, Spectrum-Based Localization Enhanced by Stack Traces (SBEST), which integrates stack trace analysis within the SBFL framework. Our method significantly improves the accuracy of fault localization by utilizing the comprehensive code execution information allied with the exception stack traces. This integration enables a more detailed understanding of system behaviour at the moment of failure, thereby addressing the gaps left by the unavailability of failing tests.
@@
  • RQ1: Are the test failures related to the bug in crash reports? We study the presence of fault-triggering tests in crash reports (i.e., bug reports that contain stack traces) and their impact on the SBFL results. We find that only 3.33% of the crash report bugs contain fault-triggering tests, which profoundly degrades the SBFL efficiency in locating these bugs. The limited execution data underscores the importance of incorporating stack traces into the SBFL process, enhancing the potential to accurately identify bug locations in the absence of traditional test failures.
@@
  • RQ2: What is the relationship between stack traces and buggy location? Given that the stack traces are the only execution information available in most of the studied bug reports, we aim to study how they relate to the buggy location and how they can be used to locate these bugs. To do so, we inspect the type of modification performed in the bugfix commit and the distance between the stack trace and the buggy methods. We find that in 98.3% of the studied bugs, the bugfix intention is directly correlated with the exception in the stack trace (e.g., adding code to handle the exception). In addition, 78.3% of the buggy methods are directly reachable from the stack traces, with an average distance of 0.34 method calls. This indicates that leveraging the stack traces to reconstruct the execution path at the time of the exception can be an effective method for pinpointing the bugs.
  • RQ3: Can we utilize the stack traces to help detect the buggy locations? Given the previous findings, we propose a new approach, SBEST, that integrates the stack trace information into Spectrum-Based Fault Localization. We compare our approach with two baselines: the Ochiai results and the position of the methods in the stack trace. Ochiai proves to be the least effective technique by a significant margin. On the other hand, we find that the stack traces alone can locate 34 out of the 60 bugs in the Top-5. This reinforces that the stack traces are a vital source of information about the bug location and should be prioritized in FL. In addition, SBEST achieves an improvement of 32.22% on MAP and 17.43% on MRR compared to the Stack Trace ranking. The enhancement of metrics suggests that the incorporation of coverage information effectively improved the ranking of stack traces, thereby providing the tool with a more comprehensive view of system execution at the time of failure.
@@
In summary, this paper presents a comprehensive study that evaluates the effectiveness of combining stack traces with test coverage data. Our findings indicate that by incorporating stack traces into the SBFL model, we can significantly improve bug localization, thereby addressing one of the key challenges in automated debugging processes. We make the data publically available online for replication and encourage future research in this area (sbe, 2024).
@@
Paper Organization. Section 2 details the background and the motivation behind this study. Section 3 surveys the related work. Section 4 explains the experimental setup. Section 5 presents the results. Section 6 discusses the implications of our findings. Section 7 presents the threats to validity. Section 8 concludes the paper and discusses future work.
2 Background and Motivation
Bug Reports contain important information for developers to fix the issues that users or other developers encounter in the software (Chen et al., 2021a). A bug report typically includes a title, a description, attached files, and comments. For example, Figure 1 shows a bug report from the Defects4j 2.0 dataset. The title offers a concise one-line description of the problem. The description provides developers with the necessary details to investigate the bug, such as the stack traces illustrating the system execution details. When additional information is required, users and developers use the comments and attached files sections to further communicate on the bug. In Figure 1 bug COMPRESS-181, a problem occurs when the system reads TAR files with symlinks. In particular, the description provides developers with a stack trace that illustrates an IOException is being thrown. The description also provides the affected nightly builds, which show no test failures.
Figure 1: An example of a bug report (COMPRESS-181). The bug report addresses a problem in reading TAR files that contain symlinks, which cause an IOException to be thrown. The report contains the stack trace of the error (gray box), and an example of TAR file (simple-aix-native-tar.tar) that triggers the bug is attached. When this bug report was created, there were no failing tests.
To assist developers in debugging such bugs, prior studies have proposed Spectrum-Based Fault Localization (SBFL) techniques (Chen et al., 2022; Lou et al., 2021; Baudry et al., 2006). SBFL relies on the information from failing tests to identify potentially faulty code. Hence, SBFL assumes that test failures that are relevant to the bug exist. SBFL works by comparing the code execution of passing and failing tests to identify their differences. The potential problematic code is then identified based on these differences. SBFL is widely studied in prior studies for its simplicity and efficiency (Santelices et al., 2009; Pearson et al., 2016). However, when there is no failing test, SBFL can not distinguish buggy code elements. For instance, in the case of bug COMPRESS-181 (Figure 1), in which there are no failing tests, SBFL can not help prioritize the faulty locations. This is because, in the absence of failing tests, there is no coverage spectrum between passing and failing executions. Thus, it becomes impossible to prioritize any code element for investigation, as SBFL relies on failing executions as a guide to identify where faults might be present. In this case, all the code locations receive the exact same suspiciousness score of zero, and the bug cannot be located.
In this paper, we investigate in detail the characteristics of Defects4j 2.0 crash reports (i.e., bug reports that contain stack traces). We dissect information such as the test failures, coverage, and stack traces by conducting a quantitative and qualitative analysis of the data and propose a new approach to address the limitations found in SBFL in such scenarios.
@@ 3.1 Fault Localization
Automated fault localization methods are a recurrent theme in the literature, being Spectrum-Based Fault Localization (SBFL) and Mutation Based Fault Localization (MBFL) two of the most famous representatives. SBFL is a very consolidated class of Fault Localization techniques. These approaches use the test information to localize the bugs, and they require at least one failing test to work. Ochiai (Abreu et al., 2007) is one of the most famous SBFL techniques. Other representatives of this class are Tarantula (Jones and Harrold, 2005), Op2 (Naish et al., 2011), BARINEL (Abreu et al., 2009) and DStar (Wong et al., 2013): similar approaches based on different formulas to compute the suspiciousness score. Given that SBFL techniques are light weighted and fast, they are also applied to other kinds of techniques such as interactive fault-localization (Gong et al., 2012) and program repair (Ye et al., 2022; Le Goues et al., 2012).
@@
MBFL techniques (Papadakis and Le Traon, 2015; Moon et al., 2014; Hong et al., 2017; Papadakis and Le Traon, 2012; Wang et al., 2022) are based on mutation analysis. For each part of the code covered by failed test cases, a set of mutants is created, and the test suite is executed. Based on the test results, the suspiciousness scores are calculated, and the most suspicious locations are ranked. Despite presenting promising results, this kind of technique is very costly and time-consuming since it requires multiple executions of the test suite. Strategies for reducing this overhead and improving the general accuracy have been studied, such as ways prioritizing the mutants (Liu et al., 2018), prioritizing the tests executed for each mutant (de Oliveira et al., 2018) and a mix of spectrum and mutant-based approaches (Cui et al., 2020). One common point between SBFL and MBFL approaches is that both types of techniques require failing tests to work. However, this might be not always the case in real life.
@@
Most of the mentioned studies use Defects4j to evaluate the results. Defects4j is one of the most famous benchmarks for test-based fault localization research. Nonetheless, recent studies (Kabadi et al., 2023; Chen et al., 2023) show that often in Defects4j, tests that were created after the bug-report creation are added to the buggy patch and thus contain developer knowledge about the bug investigation process. These tests were artificially appended to the buggy commit in Defects4j to simulate the scenario of a bug with test failures. This fact indicates that many fault localization techniques that were evaluated in Defects4j might have very different results when applied to real-world scenarios.
@@ 3.2 Bug Reports and Stack Traces
The bug report quality is a common topic in previous research. Knowing which parts of the report are more important in the bug-fixing process is very important to guide new studies on fault localization. Our focus on the stack trace information is supported by some previous results.
In the work by Bettenburg et al. (2008), the authors conduct a survey to understand what makes a good bug report from the point of view of the developers. They found that information about the steps to reproduce the bug, stack traces, and test cases are considered very helpful. In the work by Schroter et al. (2010), the authors analyze the intersection between the methods in the stack trace and the methods modified in the bugfix commit. They found out that more than 47% of the stack traces extracted from bug reports contained at least one buggy method.
An empirical study from Chen et al. (2021a) analyzes the logs’ (stack traces and log snippets) importance in crash reports. By extensively analyzing 1,561 bug reports with logs, they concluded that logs often provide helpful information about where the bug is located. They found out that in 73% of the bug reports, there is an overlap between the classes in the logs and the fixed classes.
4 Data Collection and Case Study Setup

### v2

@@ Abstract
Fault localization is a critical step in software maintenance. Yet, many existing techniques, such as Spectrum-Based Fault Localization (SBFL), rely heavily on the availability of fault-triggering tests to be effective. In practice, especially for crash-related bugs, such tests are frequently unavailable. Meanwhile, bug reports containing stack traces often serve as the only available evidence of runtime failures and provide valuable context for debugging. This study investigates the feasibility of using stack traces from crash reports as proxies for fault-triggering tests in SBFL. Our empirical analysis of 60 crash-report bugs in Defects4J reveals that only 3.33% of these bugs have fault-triggering tests available at the time of the bug report creation. However, 98.3% of bug fixes directly address the exception observed in the stack trace, and 78.3% of buggy methods are reachable within an average of 0.34 method calls from the stack trace. These findings underscore the diagnostic value of stack traces in the absence of failing tests. Motivated by these findings, we propose SBEST, a novel approach that integrates stack trace information with test coverage data to perform fault localization when fault-triggering tests are missing. SBEST shows an improvement, with a 32.22% increase in Mean Average Precision (MAP) and a 17.43% increase in Mean Reciprocal Rank (MRR) compared to baseline approaches under the scenario where fault-triggering tests are absent.
@@ 1 Introduction
Locating and fixing software bugs is an essential but often expensive and time-consuming task. Bug-fixing activities delay project progress and increase overall maintenance costs. As modern software systems grow in size and complexity, efficiently and accurately localizing faults has become increasingly critical. To address this challenge, previous research has proposed numerous different automated fault localization techniques (Abreu et al., 2007; Jones and Harrold, 2005; Naish et al., 2011; Abreu et al., 2009; Wong et al., 2013; Papadakis and Le Traon, 2015; Hong et al., 2017; Papadakis and Le Traon, 2012; Wang et al., 2022; Cui et al., 2020). Among these, Spectrum-Based Fault Localization (SBFL) stands out due to its conceptual simplicity, low computational cost, and exemplary performance in many scenarios (de Souza et al., 2016). SBFL techniques leverage information derived from test executions, specifically the spectrum of code elements executed during passing and failing tests, to calculate the probability of every code element being faulty (i.e., suspiciousness score).
Although SBFL is no longer considered state-of-the-art in terms of absolute accuracy, it remains highly relevant and widely adopted today. For example, prior learning-based fault localization techniques (Li et al., 2019; Sohn and Yoo, 2017) use SBFL scores as core input features for model training. More recently, LLM-based approaches like AutoCodeRover (Zhang et al., 2024) and LLAMAO (Yang et al., 2024) also rely on SBFL outputs to guide code analysis and method ranking. Beyond fault localization, SBFL plays a central role in various downstream software engineering tasks, including test case prioritization (Li et al., 2007; Srivastava, 2008), and automated program repair (APR) (Motwani and Brun, 2023; Kuma et al., 2020; Lou et al., 2020).
The effectiveness of SBFL highly depends on the availability of failing test cases that uncover the bug, known as fault-triggering tests (Chen et al., 2022). A fault-triggering test is a test case that fails specifically because it executes the buggy code responsible for the fault, and passes once the bug is fixed. However, studies indicate that in real-world scenarios, especially in continuous deployment and integration environments, such tests may not always be present or identified when bugs are reported (Rafi et al., 2025; Kabadi et al., 2023). The absence of fault-triggering tests can significantly impact the applicability of SBFL, as it relies on them to distinguish between faulty and non-faulty code.
In such situations, developers often rely on Crash Reports, a specific type of bug report that documents unexpected program terminations along with diagnostic details such as exception messages and stack traces. A crash report typically includes a title, description, and the stack trace that captures the program’s execution path leading to the failure (Chen et al., 2021a). For example, Figure 1 shows a crash report from the Defects4J 2.0 dataset (COMPRESS-181), which describes an IOException triggered when reading TAR files with symbolic links. The report includes the relevant stack trace but no fault-triggering tests at the time, making SBFL unable to prioritize the faulty locations. This example highlights how stack trace sequences of method calls preceding an exception can serve as an alternative source of execution information when traditional test-based approaches are not applicable (Schroter et al., 2010; Wu et al., 2014).
Figure 1: Example bug report (COMPRESS-181). The bug report addresses a problem in reading TAR files that contain symlinks, which cause an IOException to be thrown. The report contains the stack trace of the error (gray box), and an example of TAR file (simple-aix-native-tar.tar) that triggers the bug is attached. When this bug report was created, there were no failing tests.
@@
In this paper, we address the challenges of fault localization in scenarios where traditional fault-triggering test data is scarce. We conduct our study on the Defects4J 2.0 dataset (Just et al., 2014b), a widely recognized benchmark for research in fault localization and program repair. Prior studies (Kabadi et al., 2023; Rafi et al., 2025) report that some Defects4J tests were added post hoc, potentially biasing localization. To avoid this, we analyze the commit preceding the bug report to better reflect the state of the code repository that developers saw at the time. Our initial analysis of the bugs reveals that only 3.33% of crash reports, which are bug reports that contain stack traces, possess fault-triggering tests. Even for bug reports without stack traces, only around 10% of the bugs have fault-triggering tests. This low incidence of failing tests challenges the efficacy of Spectrum-Based Fault Localization (SBFL) methods, which rely heavily on test outcomes to identify bug locations. To overcome this limitation, we propose leveraging stack traces, which are rich in contextual and execution information about the bug’s root causes, as a substitute for the failing tests. We introduce a novel approach, Spectrum-Based Localization Enhanced by Stack Traces (SBEST), which integrates stack trace analysis within the SBFL framework. Our method improves fault localization accuracy in scenarios where fault-triggering tests are unavailable by combining coverage information with the execution context provided by exception stack traces. This integration enables a more detailed understanding of system behaviour at the moment of failure, thereby addressing the gaps left by the unavailability of failing tests.
@@
  • RQ1: Are the test failures related to the bug in crash reports? We examine whether the test failures observed at the bug report commit are truly fault-triggering, that is, caused by the bug described in the crash report. Our analysis shows that only 3.33% of crash-report bugs have such fault-triggering tests, limiting the applicability of SBFL. This finding underscores the need to augment SBFL with alternative sources of execution information, such as stack traces, especially when traditional failing tests are unavailable.
@@
  • RQ2: What is the relationship between stack traces and buggy location? Given that the stack traces are the only execution information available in most of the studied bug reports, we aim to study how they relate to the buggy location and how they can be used to locate these bugs. To do so, we inspect the type of modification performed in the bug fix commit and the distance between the stack trace and the buggy methods. We find that in 98.3% of the studied bugs, the bugfix intention is directly correlated with the exception in the stack trace (e.g., adding code to handle the exception). In addition, 78.3% of the buggy methods are directly reachable from the stack traces, with an average distance of 0.34 method calls. This indicates that leveraging the stack traces to reconstruct the execution path at the time of the exception can be an effective method for pinpointing bugs.
  • RQ3: Can we utilize the stack traces to help detect the buggy locations? We propose SBEST, a technique that leverages stack traces as a proxy for fault-triggering tests within the SBFL framework. To evaluate SBEST, we compare it against four baselines: Ochiai, Stack Trace ranking, SB_score (stack trace + Ochiai), and Call Frequency-Based Fault Localization. Among the baselines, Stack Trace alone ranks 34 out of 60 bugs in the Top-5, while Ochiai identifies only 2. SBEST localizes 33 bugs in the Top-5, with MAP and MRR values 32.2% and 17.4% higher than the Stack Trace baseline, respectively.
@@
In summary, this paper presents a comprehensive study evaluating the effectiveness of combining stack traces with test coverage data. Our findings indicate that incorporating stack traces into the SBFL model can enhance bug localization in scenarios where fault-triggering tests are unavailable, addressing an important challenge in automated debugging. We make the data publicly available online for replication and encourage future research in this area (Rafi et al., 2024).
@@
Paper Organization. Section 2 surveys the related work. Section 3 explains the experimental setup. Section 4 presents the results. Section 5 discusses the implications of our findings. Section 6 presents the threats to validity. Section 7 concludes the paper and discusses future work.
@@ 2.1 Fault Localization
Automated fault localization methods are a recurrent theme in the literature, being Spectrum-Based Fault Localization (SBFL) and Mutation-Based Fault Localization (MBFL) being two of the most famous representatives. SBFL is a very consolidated class of Fault Localization techniques. These approaches utilize test information to localize bugs, and they require at least one failing test to be effective. Ochiai (Abreu et al., 2007) is one of the most famous SBFL techniques. Other representatives of this class are Tarantula (Jones and Harrold, 2005), Op2 (Naish et al., 2011), BARINEL (Abreu et al., 2009), and DStar (Wong et al., 2013): similar approaches based on different formulas to compute the suspiciousness score. Given that SBFL techniques are lightweight and fast, they are also applied to other types of methods, such as interactive fault localization (Gong et al., 2012) and program repair (Ye et al., 2022; Le Goues et al., 2012).
Information Retrieval-based Fault Localization (IRFL) approaches aim to localize buggy code by treating bug reports as textual queries and comparing them against the source code using similarity metrics. Early techniques, such as BugLocator (Zhou et al., 2012), enhanced traditional IR techniques by combining TF-IDF-based scoring with historical bug-fix data to improve the ranking of relevant files. Subsequent work, such as BLUiR (Saha et al., 2013), further advanced this by leveraging structured information from source code, including method and class names, to improve retrieval accuracy over treating code as flat text. In addition, BRTracer (Wong et al., 2014) improved IR-based fault localization by segmenting source files and boosting files referenced in stack traces. Additionally, hybrid approaches, such as the one proposed by Lam et al. (2017), combine deep learning (e.g., CNNs) with IR signals to better capture the semantic relationships between bug reports and source code, outperforming purely IR-based baselines. Supervised IR-based models such as STMLocator+ (Wang et al., 2020) further enhanced accuracy by integrating textual and semantic similarity, metadata, and stack trace information.
@@
MBFL techniques (Papadakis and Le Traon, 2015; Moon et al., 2014; Hong et al., 2017; Papadakis and Le Traon, 2012; Wang et al., 2022) are based on mutation analysis. For each part of the code covered by failed test cases, a set of mutants is created, and the test suite is executed. Based on the test results, the suspiciousness scores are calculated, and the most suspicious locations are ranked. Despite presenting promising results, this kind of technique is very costly and time-consuming since it requires multiple executions of the test suite. Strategies for reducing this overhead and improving the general accuracy have been studied, such as ways of prioritizing the mutants (Liu et al., 2018), prioritizing the tests executed for each mutant (de Oliveira et al., 2018), and a mix of spectrum and mutant-based approaches (Cui et al., 2020). One common point between SBFL and MBFL approaches is that both types of techniques require failing tests to work. However, this might not always be the case in real life.
Call Frequency-Based Fault Localization (CFBFL) (Vancsics et al., 2021) is another dynamic approach that does not rely on coverage spectra but instead ranks methods based on how frequently they are executed across tests. Unlike simple count-based methods, CFBFL considers the number of times a method appears in unique call stack contexts during failing executions, filtering out noise from loops and repetitive calls. A related approach by  Beszedes et al. (2020) improves SBFL by analyzing function call chains, computing suspiciousness over call paths, and then attributing it to individual functions. Both techniques enhance fault localization by leveraging richer execution context beyond binary coverage.
@@
Most of the mentioned studies use Defects4J to evaluate the results. Defects4J is one of the most widely used benchmarks for fault localization techniques that rely on test coverage to identify suspicious code elements. Nonetheless, recent studies (Kabadi et al., 2023; Rafi et al., 2025) show that often in Defects4J, tests created after the bug report are added to the buggy patch and thus contain developer knowledge about the bug investigation process. These tests were artificially appended to the buggy commit in Defects4J to simulate a scenario with a bug and test failures. This fact suggests that many fault localization techniques evaluated in Defects4J may yield significantly different results when applied to real-world scenarios.
@@ 2.2 Bug Reports and Stack Traces
Bug report quality has long been a central topic in software engineering research. Understanding which elements of a bug report are most helpful in identifying and resolving faults can help guide the development of more effective fault localization techniques. One of the most consistently valuable components of bug reports is the stack trace, particularly in the context of crashes and runtime errors.
Bettenburg et al. (2008) conducted a developer survey to determine what makes a bug a helpful report. Their results highlight the importance of detailed reproduction steps, stack traces, and associated test cases. These elements help developers quickly understand and reproduce the issue, making stack traces especially valuable for debugging. Schroter et al. (2010) further investigated the connection between stack traces and the actual faults by analyzing whether methods in the reported stack traces were modified during the corresponding bug fix. Their findings revealed that in over 47% of the cases, at least one buggy method appeared in the stack trace, suggesting that stack traces often point directly to fault-relevant areas of the code.
Building on this, Chen et al. (2021a) performed a large-scale empirical study of 1,561 bug reports containing logs, including stack traces and log messages. They found that in 73% of the cases, there was an overlap between the classes mentioned in the logs and the classes involved in the bug fix. This reinforces the view that logs and especially stack traces provide strong signals for localizing faults. Several techniques have been proposed that leverage stack traces more directly for automated fault localization. For example, Jiang et al. (2012) use stack traces in conjunction with program slicing to localize null pointer exceptions better. Gong et al. (2014) generates stack traces to aid in crash fault localization and demonstrates their approach’s effectiveness by identifying 64% of crashing faults in Firefox 3.6.
CrashLocator (Wu et al., 2014) introduced a scalable method for leveraging stack traces across multiple crash reports. It clusters similar crash reports into buckets corresponding to the same fault and cross-analyzes the slightly varying stack traces within each bucket. The approach expands stack traces using static call graph traversal and ranks methods based on their frequency within the cluster, rarity across other clusters, and proximity to the reported crash point. This design allows CrashLocator to combine statistical signals with structural code context, yielding highly accurate fault predictions even in large-scale systems. Together, these studies and tools highlight the value of stack traces not only as debugging aids for developers but also as key artifacts that can be mined and augmented for automated fault localization.
Comparison with Closely Related Studies. While prior studies have demonstrated the usefulness of stack traces for fault localization, their objectives and assumptions differ from ours. CrashLocator (Wu et al., 2014) clusters multiple crash reports and expands stack traces using static call graphs to rank methods by their frequency and proximity within crash clusters. In contrast, SBEST focuses on individual crash reports and integrates stack-trace-derived execution paths directly into the Spectrum-Based Fault Localization (SBFL) framework, treating stack-trace methods as proxies for fault-triggering tests. Call Frequency-Based Fault Localization (CFBFL) (Vancsics et al., 2021) extends Spectrum-Based Fault Localization by weighting coverage information with the frequency of method occurrences in stack traces collected during test executions. Information Retrieval (IR)-based techniques such as BugLocator (Zhou et al., 2012) and STMLocator+ (Wang et al., 2020) rely mainly on textual similarity between bug reports and source code, occasionally incorporating stack traces as supplementary input. In contrast, SBEST leverages stack traces extracted from crash reports rather than from executed tests and integrates them with coverage data to perform fault localization even when fault-triggering tests are missing.
3 Data Collection and Study Design